What volume of a 0.171 M barium hydroxide solution is required to neutralize 24.0 mL of a 0.225 M hydrobromic acid solution?

_________ mL barium hydroxide
1
Expert's answer
2019-03-18T02:22:41-0400

 xmol    0.0054mol

Ba(OH)2 + 2HBr = BaBr2 +2H2O

 1mol      2mol

 

n(HBr) = C(HBr) × V(HBr) = 0.225M × 0.024L = 0.0054mol.

n(Ba(OH)2) = x = 0.0054 × 1 / 2 mol = 0.0027mol.

V(Ba(OH)2) = n(Ba(OH)2) / C(Ba(OH)2) = 0.0027mol / 0.171M = 0.0158L = 15.8mL.